# Semi-Supervised Learning for Lensed Quasar Detection

### v1

@@ 3.5 Virtual Adversarial Training
VAT is a technique designed to enhance the robustness and accuracy of deep learning models in tasks where labelled data are scarce but unlabelled data are abundant. Readers are referred to Miyato et al. (2017) for full details, but we give a brief overview here. At its core, VAT operates by augmenting the labelled data by classifying the unlabelled data and then penalising the model if the classifications change after a small adversarial perturbation to the original image. To achieve this, the loss function is composed of two terms. The first is the normal loss associated with the chosen deep classifier, in our case cross entropy loss, together with any regularisation. The second term is calculated by classifying the unlabelled images, making a small adversarial perturbation and then classifying the perturbed images. The number of images, multiplied by a tuneable hyperparameter, which change their classification under this perturbation becomes the second loss term for the model. Adversarially perturbing the image with a deep model is straightforward as the derivative of the image can be taken with respect to the final classification to give a vector pointing in the adversarial direction. The image is then modified by adding this vector to the image to create a new image which has been modified in the “worst” possible way for the classifier. This new image can then simply be passed through the model to result in another classification to be compared to the old classification.
@@
In this work we train a convolutional neural network using VAT. We structure it similarly to the autoencoder described in Section 3.2, starting three channels in the input layer and then having four convolutional layers with 8, 16, 32, and 64 channels. Each layer has a kernel size of 3, a stride of 2 and a padding of 1. Between each convolutional layer was a batch normalisation (Ioffe & Szegedy, 2015) layer and a leaky ReLU activation function. We found that the best F1 score was achieved by flattening the network to 260 nodes before passing through 4 densely connected layers which have 256, then 128, 64 and 2 features. Between each of the densely connected layers is a leaky ReLU activation function and a batch normalisation layer, with a sigmoid activation after the final densely connected layer. We experimented with varying values of dropout to regularise the network, but ultimately found that the network performed best without dropout.
@@ 4.1 Comparison
Figure 8: Some example classifications of lensed quasars from the best autoencoder-classifier model. The top row are two correctly classified lensed quasars (as indicated by the green border), while the bottom row are two lensed quasars which the model fails to identify (as indicated by the red border). In each image the black dashed square indicates the inner 32x32 pixels (8x8 arcseconds) which were passed to the autoencoder. In the bottom left image the model likely misclassifies the lens as it has an extremely wide separation resulting in the some of the quasar images being cropped out. In the bottom right image the model may conflate the structure with a single source bearing some visual artefacts.
The best performing autoencoder-classifier model achieved an F1 score of 0.87 on the test set, significantly outperforming the best VAT model with an F1 score of 0.58. Thus, on our relatively clean data set the simplicity of the traditional classifier outperforms the flexibility of the convolutional neural network trained via VAT. Some example classifications from the autoencoder-classifier model are shown in Figure 8.
@@ 4.2 On-sky Results
GRALJ121709.65-025621.60 2023 Jan 22 122 2$\times$300 $z=1.464$ quasar + $z=0.165$ emission-line galaxy
GRALJ140833.73+042229.98 2023 Jan 22 173 2$\times$600 $z=2.998$ quasar lensed by $z=0.542$ early-type galaxy
@@ 4.3 Model Improvements
Although significantly increasing the number of confirmed lensed quasars in the training data is infeasible using the approach described in this work, progress could be achieved by incorporating simulated images into the training data. This was done by Akhazhanov et al. (2022), which produced significantly larger labelled sets. However, it introduces challenges with corrupting the simulated images to match the Pan-STARRS/DESI images. Another difficulty is that by training on lensed quasars from simulations, we bias the model toward detecting objects like those that we generate, so care has to be taken to generate a sufficiently diverse range of lensed quasar simulations.
@@ 5 Conclusions
Despite these challenges, we develop two models which are capable of distinguishing between lensed and unlensed quasars from image data. The first approach is to train an $\beta$-VAE to recreate images of quasars, of which we source a few million. The labelled data is passed through the $\beta$-VAE and the latent space is passed to a classifier, the best performing of which was found to be gradient boosting. The second approach is to train a convolutional neural network using VAT to allow the network to be trained on both labelled and unlabelled data, thus taking advantage of the significant number of quasars to increase the performance of the network despite our small set of labelled data. Of these two approaches we find that the autoencoder-classifier model achieves a better F1 score, 0.87, on our clean dataset compared to our VAT model, 0.58. However, when the models are tasked with finding undiscovered lensed quasars they perform with roughly equal success as the end-to-end approach of the VAT model allows it to generalise well. Both models still struggle with images which have few similarities to images in the labelled set (e.g., crowded fields). We conducted observations at the W.M. Keck Observatory of 5 lensed quasar candidates delivered by our classifiers, confirming one to be a previously undiscovered lensed quasar GRALJ140833.73+042229.98, three others to be interlopers consisting of a quasar and a nearby star and one candidate an unconfirmed lens requiring further follow-up.
@@ Acknowledgements
In addition to those mentioned in the main text, we would also like to acknowledge other python packages used as part of this work: NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007), seaborn (Waskom, 2021), pandas (Wes McKinney, 2010; pandas development team, 2020), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), astroquery (Ginsburg et al., 2019), PyTorch (Paszke et al., 2019), scikit-learn (Pedregosa et al., 2011), Pillow (Umesh, 2012), scikit-image (van der Walt et al., 2014), VAT-pytorch (Lyakaap, 2018) and ray (Moritz et al., 2017).
@@
The code underlying this article are available on Zenodo (Sweeney, 2024). The data may be made available upon personal request.
Appendix A Tuning the classifiers
@@
The code used to tune the models is available on Zenodo (Sweeney, 2024).

### v2

@@ 3.5 Virtual Adversarial Training
VAT is a technique designed to enhance the robustness and accuracy of deep learning models in tasks where labelled data are scarce but unlabelled data are abundant. Readers are referred to Miyato et al. (2017) for full details, but we give a brief overview here. At its core, VAT operates by augmenting the labelled data by classifying the unlabelled data and then penalising the model if the classifications change after a small adversarial perturbation to the original image. To achieve this, the loss function is composed of two terms. The first is the normal loss associated with the chosen deep classifier, in our case cross entropy loss, together with any regularisation. The second term is calculated by classifying the unlabelled images, making a small adversarial perturbation and then classifying the perturbed images. The number of images, multiplied by a tuneable hyperparameter, which change their classification under this perturbation becomes the second loss term for the model. Adversarially perturbing the image with a deep model is straightforward as the derivative of the image can be taken with respect to the final classification to give a vector pointing in the adversarial direction. The image is then modified by adding this vector to the image to create a new image which has been modified in the worst possible way for the classifier. This new image can then simply be passed through the model to result in another classification to be compared to the old classification.
@@
In this work we train a convolutional neural network using VAT. We structure it similarly to the autoencoder described in Section 3.2, starting three channels in the input layer and then having four convolutional layers with 8, 16, 32, and 64 channels. Each layer has a kernel size of 3, a stride of 2 and a padding of 1. Between each convolutional layer was a batch normalisation (Ioffe and Szegedy, 2015) layer and a leaky ReLU activation function. We found that the best F1 score was achieved by flattening the network to 260 nodes before passing through 4 densely connected layers which have 256, then 128, 64 and 2 features. Between each of the densely connected layers is a leaky ReLU activation function and a batch normalisation layer, with a sigmoid activation after the final densely connected layer. We experimented with varying values of dropout to regularise the network, but ultimately found that the network performed best without dropout.
@@ 4.1 Comparison
Figure 8: Some example classifications of lensed quasars from the best autoencoder-classifier model. The top row are two correctly classified lensed quasars (as indicated by the green border), while the bottom row are two lensed quasars which the model fails to identify (as indicated by the red border). In each image the white dashed square indicates the inner 32x32 pixels (8x8 arcseconds) which were passed to the autoencoder. In the bottom left image the model likely misclassifies the lens as the reddening caused by the lensing galaxy affects one image significantly more than the other. In the bottom right image the model may be confused by the colouring induced by the $i$ band (blue channel) being null in this area of the sky in the DESI survey.
The best performing autoencoder-classifier model achieved an F1 score of 0.897 on the test set, significantly outperforming the best VAT model with an F1 score of 0.58. Thus, on our relatively clean data set the simplicity of the traditional classifier outperforms the flexibility of the convolutional neural network trained via VAT. Some example classifications from the autoencoder-classifier model are shown in Figure 8.
@@ 4.2 On-sky Results
[…]
@@ 4.3 Model Improvements
Although significantly increasing the number of confirmed lensed quasars in the training data is infeasible using the approach described in this work, progress could be achieved by incorporating simulated images into the training data. This was done by Akhazhanov et al. (2022), which produced significantly larger labelled sets. However, it introduces challenges with corrupting the simulated images to match the Pan-STARRS/DESI survey images. Another difficulty is that by training on lensed quasars from simulations, we bias the model toward detecting objects like those that we generate, so care has to be taken to generate a sufficiently diverse range of lensed quasar simulations.
@@ 5 Conclusions
Despite these challenges, we develop two models which are capable of distinguishing between lensed and unlensed quasars from image data. The first approach is to train an $\beta$-VAE to recreate images of quasars, of which we source a few million. The labelled data is passed through the $\beta$-VAE and the latent space is passed to a classifier, the best performing of which was found to be a densely connected artificial neural networkgradient boosting. The second approach is to train a convolutional neural network using VAT to allow the network to be trained on both labelled and unlabelled data, thus taking advantage of the significant number of quasars to increase the performance of the network despite our small set of labelled data. Of these two approaches we find that the autoencoder-classifier model achieves a better F1 score, 0.897, on our clean dataset compared to our VAT model, 0.58. However, when the models are tasked with finding undiscovered lensed quasars they perform with roughly equal success as the end-to-end approach of the VAT model allows it to generalise well. Both models still struggle with images which have few similarities to images in the labelled set (e.g., crowded fields). We conducted observations at the W.M. Keck Observatory of 5 lensed quasar candidates delivered by our classifiers, confirming one to be a previously undiscovered lensed quasar GRALJ140833.73+042229.98, three others to be interlopers consisting of a quasar and a nearby star and one candidate an unconfirmed lens requiring further follow-up.
@@ Acknowledgements
In addition to those mentioned in the main text, we would also like to acknowledge other python packages used as part of this work: NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007), seaborn (Waskom, 2021), pandas (McKinney, 2010; pandas development team, 2020), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), astroquery (Ginsburg et al., 2019), PyTorch (Paszke et al., 2019), scikit-learn (Pedregosa et al., 2011), Pillow (Umesh, 2012), scikit-image (van der Walt et al., 2014), VAT-pytorch (Lyakaap, 2018) and ray (Moritz et al., 2017).
@@
The code underlying this article are available on Zenodo (Sweeney, 2025). The data may be made available upon personal request.
Appendix A Pan-STARRS1 JPEG Compression
The Pan-STARRS1 cutout images are generated via a two-step process, here I paraphrase the relevant section of an email from Rick White (priv. comm., 2025) from the Pan-STARRS1 data processing team describing this process:
  1. 1.
    “The image contrast is limited using a linear stretch between the 0.5% and 99.5% pixel values within the cutout. This process clips the occasional bad pixel or bright star.
  2. 2.
    Then an arc-hyperbolic sine (asinh) transform is applied to the scaled pixels values. The asinh function behaves similarly to the log function for large values ($\gg 1$ or $\ll-1$) but is linear for values near zero. It is also (unlike log) defined for negative pixels. This behaviour enhances contrast near the sky and further suppresses the brightness of bright stars.
The underlying C code for the actual equations for the contrast computation can be found here: https://github.com/spacetelescope/fitscut
The clipping of bad pixels and asinh transformation of the pixel values likely result in the increased performance observed in this work.
Appendix B F1, Recall and Precision Scores
A precision score is defined as follows:
Precision $=\frac{\text{True positives}}{\text{True positives + False positives}}$ (2)
A recall score is defined as follows:
Recall $=\frac{\text{True positives}}{\text{True positives + False negatives}}$ (3)
An F1 score is the harmonic mean of precision and recall:
$\text{F}_{1}$ $=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision + Recall}}$ (4)
Appendix C Tuning the classifiers
@@
The artificial neural network wasn’t tuned using RandomizedSearchCV as the custom implementation didn’t easily fit into this framework. Instead, we used StratifiedKFold and a grid search over the following sets of hidden layers: (64, 64), (64, 32), (32, 32), (64, 64, 64), (64, 32, 16), (32, 32, 32), (64, 64, 64, 64), (64, 32, 16, 8) and (32, 32, 32, 32), combined with either dropout (tested values of 0.05, 0.1, 0.125, 0.15, 0.175 and 0.2) or weight decay ($10^{-4}$, $3\cdot 10^{-3}$, $10^{-3}$, $3\cdot 10^{-2}$, $10^{-2}$) with PyTorch’s BCEWithLogitsLoss as the loss function. Because of the variation model performance due to the randomised initial weights and biases of the neural network, the best combination of hidden layers and regularisation varied on consecutive runs. To find the best set of hyperparameters StratifiedKFold run 7 times with seeds from 0–6 and the best model was selected based on the following criteria:
  1. 1.
    The set of hidden layers which appeared the most times as the best set of hidden layers.
  2. 2.
    If the regularisation differs amongst solutions, the value which appears most frequently is chosen.
  3. 3.
    Ties are broken by the lowest average binary cross entropy error.
The code used to tune the models is available on Zenodo (Sweeney, 2025).